# Extension of the integrated HydroKinetic Model to BES RHIC and GSI-FAIR nuclear collision energies

### v2

@@ Abstract
The present work is devoted to developing the integrated hydrokinetic approach (iHKM) for relativistic nucleus-nucleus collisions. While the previous cycle of works on this topic focused on ultra-relativistic collisions at the top RHIC and different LHC energies, the current work addresses relativistic collisions at the lower energies, specifically ranging from approximately 1 to 50 GeV per nucleon pair in the center-of-mass colliding system. In such collisions, the formation times for the initial state of dense matter can be up to three orders of magnitude longer than those in ultra-relativistic collisions. This difference reflects a fundamentally distinct nature and formation process, particularly regarding the possible stages of initial state evolution, including thermalization (which may be only partial at very low collision energies), subsequent hydrodynamic expansion, and the final transition of matter evolution into a hadronic cascade. These stages, which are fully realized in ultra-relativistic reactions, can also occur within the energy range of BES RHIC, albeit with distinct time scales. This publication not only advances the theoretical development of iHKM (referred to, if necessary, as the extended version of integrated Hydrokinetic Model, iHKMe), but also provides examples of model applications for calculating observables. A systematic description across a wide range of experimental energies, which is preliminary yet quite satisfactory, for spectra, flow, and femtoscopy, will follow this study.
@@ I Introduction
The intermediate and low relativistic range of collision energy experiments are of special interest. The currently acting ones are associated with the Beam Energy Scan program at RHIC (BES RHIC) and the HADES experiment at the GSI accelerator facility. The nearest planning is the Compressed Baryon Matter (CBM) experiment at the GSI-FAIR. Despite collision processes at the LHC, where the transition between hadron and quark-gluon matters is happening without the phase transition (crossover), at the above-mentioned experiments characterized by large net baryon densities in creating matter, one hopes to search for the thermodynamic line of the phase transition between hadron and quark-gluon matters and also try to discover the critical endpoint [14].
@@
The objective of this paper is to present an extended integrated hydrokinetic model for the soft physics in all mentioned experiments, covering (all together with already developed iHKM for ultra-relativistic energies) the range from 1 GeV up to 10 TeV energies per nucleon colliding pair in their center of mass, within a unified approach based on the extended integrated HydroKinetic Model - iHKMe. The latter will supplement the iHKM, which is already available for ultra-relativistic energies, to the intermediate and low relativistic collision energies. For each considered energy, all the possible stages of nuclei collision processes will be investigated within the same unified description as in the original iHKM. The mentioned stages of the matter evolution during the collision process are: the formation of the initial conditions for the system expansion into a vacuum just after the collision, gradual thermalization of created superdense matter (maybe not complete for quite low collision energies, that can happen without hydro-stage), its consequent hydrodynamic evolution for intermediate energies, the translation of the description of hydrodynamic/or partially thermal medium to particle language – so-called particlization, and, finally, the cascade stage for still interacting already individual particles [15]. So, our approach also includes the possibility that at fairly low energies, not all of these stages (e.g., hydro-stage) will be realized/activated.
@@
Our idea, and later its full realization, was: to get more intensive flow, both the radial and anisotropic, one does not need the pressure gradient in the created fireball and, so, does not need the early fast thermalization [24]. These flows can be well developed because only the geometrical form of the very initial system, which is essentially finite in nucleus-nucleus collisions and, in addition, has an anisotropic shape in non-central collisions. What is very important - in any case, one needs the system’s thermalization sooner or later to describe the data, but not necessarily a fast one. If it happens, say, at 1 fm/$c$, the transverse flow will be present already at this (relatively late) thermalization/ hydrodynamization time because they were developed at the pre-thermal stage, even with quite gradually appearing pressure. The corresponding formalism developed allows one to build the full model of heavy ion collisions that incorporates all stages of heavy ion collision processes without the physically controversial hypothesis of the very early system’s thermalization [24, 15].
@@
It is clear that the model of initial conditions for the consequent pre-thermal (thermalization) stage are very different. While in iHKM we use a hybrid approach based on MC-GLAUBER calculations, realized in the GLISSANDO-2 model [25] in the transverse direction, and Color Glass Condensate in the longitudinal one, the initial conditions are dramatically changed in the collisions at BES RHIC and below energies where the overlapping time is differed (larger) by 3 order of the value. The initial conditions in the extended for intermediate and small relativistic heavy ion collision energies model - iHKMe are based on the quasiclassical UrQMD simulations [26] with added quantum (and classical) fluctuations during the thermalization process. It leads to partial thermal evolution of the matter, created at small and intermediate collision energies.
The goal of this paper is to propose the theoretical basis for the description of the soft physics at the intermediate and small relativistic energies, in addition to what has been done already for ultrarelativistic energies (iHKM), and also to illustrate the results of the developed model at one of the collision energies. A consistent description of the observables within iHKMe/iHKM in BES RHIC and GSI-FAIR experiments will be presented in subsequent publications.
@@ II.1 The basic aspects of the approach
In the extended version of iHKM - we call it iHKMe - we rely on the transport UrQMD model for the initial pre-equilibrium dynamics at low and intermediate relativistic collision energies. The possibility of using for the same aim in other similar models like SMASH [27] is planned in future publications.
The problem of transformation of the hydrodynamic approach to particle evolution language and vice versa is quite non-trivial. Especially if it concerns the low enough but still relativistic energies, e.g., at $\sqrt{s}<10$A GeV for $A+A$ collisions. The reasons for this are the following.
First, the obvious problem is that during UrQMD (or SMASH) simulations of the initial stage of collisions, the gradual transformation, at least of part of the total system, to a hydrodynamic subsystem cannot be based on the initial singular distribution of particles in one cascade event. So, one needs to use some smoothing/averaging even for event-by-event analysis.
@@
Second, the system may not be hydrodynamized completely at very low energies. The hydrodynamically expanding part of the matter transforms into final particles just into the surrounding part of the interacting hadron gas, which escaped the hydrodynamic phase during the total matter evolution. The process of transition of the hydrodynamically expanding part of the matter into hadrons is called particlization and is the starting point for the consequent stage of hadron cascade evolution. In this picture, the main theoretical point is the explanation of the real physical mechanism of thermalization/hydrodynamization. The approach [15], which we use now, is phenomenologically described by utilizing the thermalization and relaxation times within the Boltzmann equation approximation. The gradual transformation of the initially non-thermal matter into (partially) thermalized matter is governed by conservation law equations.
@@
Third, the important question concerns the nature of the thermal fluid that appears against the backdrop of the initial UrQMD (or SMASH) gas of colliding hadrons. We suppose that initially, it appears due to local quantum and classical density fluctuations through the QCD droplet formation. Similar to our first point (see above), even at event-by-event analysis, the smoothing procedure is necessary to describe some stage of the collision process in the continuous medium approximation. We just smooth out the droplet picture for very high-density local fluctuations into an effectively hydrodynamic one.
@@
The realistic situation, however, can be more complicated, even at the selected criteria: in the case of low collision energies, the system may never reach complete thermalization, but only if possible, a partial one. In such a case, in spite of near full thermalization at the particlization stage at high enough energies, the system at the decaying (into the interacting particles) stage consists of two expanding components of baryon-rich matter: particle gas and fluid. The most natural way then is to select the energy density interval when the mechanism, forming the dense (QGP?) droplets, stops working, and soon after the corresponding (see above) hydrodynamical part of the system decays into hadrons without reaching full thermalization. In our approximation, it means that one has to transform the hydrodynamically involved part of the system into particles at this droplet’s (mean) decay time. So, because we do not know the quantum component of the QCD fluctuations, the corresponding energy density at which the decay of the hydro-component at small relativistic energies happens is a free parameter. The corresponding particle injection from this decay of the hydro-component in the case of not full thermalization is just added to the preserved yet UrQMD component. Of course, the local energy-momentum conservation law for the evolution and hydro-decays of the total system, consisting of both UrQMD (or SMASH) and hydro-components, must be implemented in any scenario. It is done in the iHKMe (like in iHKM) scenario that is developed in the presented article.
@@
However, as we already marked, switching from quasiclassical microscopic models (UrQMD, SMASH, etc,..) to the macroscopic hydrodynamic regime could be impossible if one naively tries to based on a distribution function from a single transport simulation, it brings significant fluctuations in coordinates and anisotropies in momentum space, contradicting the basic assumptions of hydrodynamics. So, one needs to use some averaging procedure to provide smooth initial conditions for a subsequent description of hydrodynamic expansion and, of course, especially, for event-by-event analysis.
@@
There are two common solutions to this problem. The straightforward one is to generate huge amounts of events and then average over them. However, to study the influence of fluctuations in the initial conditions on the final observables, one needs to introduce some similarity criteria between events. Of course, averaging over the centrality class might be too rough and unsuitable for event-by-event analysis. The common approach to this problem is applying some Gaussian smearing procedure to each particle and then constructing only the time component of the stress-energy tensor $T^{0\mu}$ and baryon current $J^{0}_{\mu}$. The other components are restored using an equation of state and explicit representation of the relativistic hydrodynamic tensors through macroscopic fields of velocity $u^{\mu}$, energy-density $\epsilon$, pressure $p$, and baryon density $n_{B}$.
@@
where ${\bf u}$ is velocity of the particle and $\gamma$ is the Lorentz contraction factor. Such a procedure can be attributed to the averaging over an ensemble of “similar” collision events without generating them. The similarity between the events is described by the $\sigma$ parameter.
@@
Here, $r_{ij}^{\mu}=x^{\mu}_{i}-x^{\mu}_{j}$ represents the radius vector between the particle and the center of the cell $\left(x_{i},x_{j}\in\sigma^{\mu}\right)$, $n_{i}^{\mu}$ is the normal vector to hypersurface at point $x_{i}$, while $R$ is a free scalar parameter. We utilize the kernel (8) on hypersurfaces of constant proper time $\tau=\text{const}$. Then, for a cell with space-rapidity $\eta_{j}$,
@@
where $\Delta x,\Delta y,\Delta\eta$ are cell sizes in three directions, with the numerical values $0.3$ fm, $0.3$ fm, and $0.05$ respectively.
Utilizing kernel (8) we obtain the following inputs for the next stages of the model:
@@
Lastly, let us note that in this paper we do not consider separate equations for electric charge and strangeness, assuming locally $n_{s}=0$ and $n_{q}=A/Zn_{B}\approx 0.4n_{B}$ for gold nuclei.
@@ II.2 Thermalization
One of the key aspects of iHKM is the thermalization stage, in other words, the hydrodynamization process. During this stage, matter can be phenomenologically split into two distinct components: a (near) local equilibrium one, described by macroscopic fields, and non-thermalized components represented by out-of-equilibrium hadrons and strings evolving via UrQMD. These components contribute to the non-equilibrium energy-momentum tensor [12] and the charge currents. so that the corresponding equations guarantee the conservation laws of the corresponding values for the total systems.
@@
where ${\cal P}_{\tau}={\cal P}(\tau)$ is a weight function such that ${\cal P}(\tau_{0})=1$ at the start of the thermalization stage, ${\cal P}(\tau_{th})=0$ at the end, and $10<{\cal P}(\tau_{0}<\tau<\tau_{th})<$ 1 in between. Its explicit form will be discussed later. Both the total stress-energy tensor and the baryon current obey conservation laws:
@@
$\partial_{\mu}T^{\mu\nu}_{\text{urqmd}}(x)=0,\quad\partial_{\mu}J_{\text{urqmd}}^{\mu}(x)=0.$ (17)
@@
$\tilde{J}^{\mu}_{\text{hydro}}(x)=T^{\mu}_{\text{hydro}}(x)\cdot\left(1-{\cal P}_{\tau}\right),$ (21)

### v3

# Extension of the integrated hydrokinetic model to nuclear collision energies relevant for the RHIC Beam-Energy Scan program and the research program at GSI-FAIR

@@ Abstract
The present work is devoted to developing the integrated hydrokinetic model (iHKM) for relativistic nucleus-nucleus collisions. While the previous cycle of works on this topic focused on ultra-relativistic collisions at the top RHIC and different LHC energies, the current work addresses relativistic collisions at the lower energies, specifically ranging from approximately 2 to 50 GeV per nucleon pair in the center-of-mass colliding system. In such collisions, the formation times for the initial state of dense matter can be up to three orders of magnitude longer than those in ultra-relativistic collisions. This difference reflects a fundamentally distinct nature and formation process, particularly regarding the possible stages of initial state evolution, including thermalization (which may be only partial at very low collision energies), subsequent hydrodynamic expansion, and the final transition of matter evolution into a hadronic cascade. These stages, which are fully realized in ultra-relativistic reactions, can also occur within the energy range of BES RHIC, albeit with distinct time scales. This publication not only advances the theoretical development of iHKM (referred to, if necessary, as the extended version of integrated Hydrokinetic Model, iHKMe), but also provides examples of model applications for calculating observables. A systematic description across a wide range of experimental energies, which is preliminary yet quite satisfactory, for spectra, flow, and femtoscopy, will follow this study.
@@ I Introduction
Experiments in the intermediate and low relativistic collision energy range are of special interest. The currently acting ones are associated with the Beam Energy Scan program at RHIC (BES RHIC) and the HADES experiment at the GSI accelerator facility. The nearest planning is the Compressed Baryon Matter (CBM) experiment at the GSI-FAIR. Despite collision processes at the LHC, where the transition between hadron and quark-gluon matters is happening without the phase transition (crossover), at the above-mentioned experiments characterized by large net baryon densities in creating matter, one hopes to search for the thermodynamic line of the phase transition between hadron and quark-gluon matters and also try to discover the critical endpoint [14].
@@
The objective of this paper is to present an extended integrated hydrokinetic model for the soft physics in all mentioned experiments, covering (all together with already developed iHKM for ultra-relativistic energies) the range from 2 GeV up to 10 TeV energies per nucleon colliding pair in their center of mass, within a unified approach based on the extended integrated HydroKinetic Model - iHKMe. The latter will supplement the iHKM, which is already available for ultra-relativistic energies, to the intermediate and low relativistic collision energies. For each considered energy, all the possible stages of nuclear collision processes will be investigated within the same unified description as in the original iHKM. The mentioned stages of the matter evolution during the collision process are: the formation of the initial conditions for the system expansion into a vacuum just after the collision, gradual thermalization of created superdense matter (maybe not complete for quite low collision energies, that can happen without hydro-stage), its consequent hydrodynamic evolution for intermediate energies, the translation of the description of hydrodynamic/or partially thermal medium to particle language – so-called particlization, and, finally, the cascade stage for still interacting already individual particles [15]. So, our approach also includes the possibility that at fairly low energies, not all of these stages (e.g., hydro-stage) will be realized/activated.
@@
Our idea, and later its full realization, was: to get more intensive flow, both radial and anisotropic, one does not need the pressure gradient in the created fireball and, so, does not need the early fast thermalization [24]. These flows can be well developed because only the geometrical form of the very initial system, which is essentially finite in nucleus-nucleus collisions and, in addition, has an anisotropic shape in non-central collisions. What is very important - in any case, one needs the system’s thermalization sooner or later to describe the data, but not necessarily a fast one. If it happens, say, at 1 fm/$c$, the transverse flow will be present already at this (relatively late) thermalization/ hydrodynamization time because they were developed at the pre-thermal stage, even with quite gradually appearing pressure. The corresponding formalism developed allows one to build the full model of heavy ion collisions that incorporates all stages of heavy ion collision processes without the physically controversial hypothesis of the very early system’s thermalization [24, 15].
@@
It is clear that the model of initial conditions for the consequent pre-thermal (thermalization) stage is very different. While in iHKM we use a hybrid approach based on MC-GLAUBER calculations, realized in the GLISSANDO-2 model [25] in the transverse direction, and Color Glass Condensate in the longitudinal one, the initial conditions are dramatically changed in the collisions at BES RHIC and below energies where the overlapping time is differed (larger) by 3 order of the value. The initial conditions in the extended for intermediate and small relativistic heavy ion collision energies model - iHKMe are based on the quasiclassical UrQMD simulations [26] with added quantum (and classical) fluctuations during the thermalization process. It leads to partial thermal evolution of the matter, created at small and intermediate collision energies.
The goal of this paper is to propose the theoretical basis for the description of the soft physics at the intermediate and small relativistic energies, in addition to what has been done already for ultrarelativistic energies (iHKM), and also to illustrate the results of the developed model at one of the collision energies. A consistent description of the observables within iHKMe/iHKM in BES RHIC, and GSI-FAIR experiments will be presented in subsequent publications.
@@ II.1 The basic aspects of the approach
In the extended version of iHKM, hereafter referred to as iHKMe, when necessary, the initial pre-equilibrium dynamics at low and intermediate relativistic collision energies are modeled using the UrQMD hadron-string cascade approach [26]. This approach was first employed as an initial stage for hydrodynamic simulations in Ref. [27]. It offers several advantages: it describes non-equilibrium dynamics and provides event-by-event fluctuating, $3+1$ dimensional distributions of energy, momentum, and conserved charge densities (baryon, electric, and strangeness), while explicitly conserving these quantities. Notably, it enables the use of the same microscopic model for both the initial and final (afterburner) stages, allowing for a continuous description of the system’s peripheral regions (the corona), where thermalization is not achieved. Our model architecture also allows the incorporation of other similar transport approaches, such as SMASH [28], JAM [29], PHSD [30], and GiBUU [31]. This includes variants with density-dependent potentials that influence the equation of state [32, 33], which may be particularly relevant at a few GeV collision energies due to the importance of the initial compression stage, as demonstrated in Ref. [34]. However, these alternatives are not currently implemented.
The transformation between the hydrodynamic description and the microscopic particle evolution presents significant challenges, particularly at relatively low but still relativistic collision energies, such as $\sqrt{s_{NN}}<10$ GeV in $A+A$ collisions. The underlying reasons for this complexity are the following. First, a fundamental difficulty arises because during UrQMD (or other transport model) simulations of the initial collision stage, the gradual formation of a hydrodynamic subsystem from the total system cannot be based on the initial singular particle distribution from a single cascade event. Therefore, some smoothing or averaging procedure is required even in event-by-event analyses (a comprehensive review of this topic is provided in Ref. [35]).
@@
Second, the system may not become fully hydrodynamized at very low collision energies. In such cases, the hydrodynamically expanding part of the matter transforms into final-state particles within a surrounding region of the interacting hadron gas that never entered the hydrodynamic phase during the system’s evolution. The process by which the hydrodynamic matter converts into hadrons is referred to as particlization, and it serves as the starting point for the subsequent hadronic cascade stage in the total system. A central theoretical challenge in this framework is to describe the physical mechanism underlying thermalization or hydrodynamization. The approach employed in Ref. [15], which is currently used in our model, provides a phenomenological description based on the Boltzmann equation with relaxation and thermalization times. The gradual transformation of initially non-thermal matter into (partially) thermalized matter is governed by conservation law equations. Other attempts to model the gradual fluidization of nonequilibrium matter have been proposed. For example, in Ref. [36], strings are dynamically formed between participant quarks or nucleons and subsequently thermalize, serving as a source term for the hydrodynamic evolution. A similar mechanism of continuous fluidization from strings or hadronic resonances has been introduced in the context of the UrQMD and JAM transport models in Refs. [37, 38], respectively.
@@
Third, the important question concerns the nature of the thermal fluid that appears against the backdrop of the initial gas of colliding hadrons. We suppose that initially, it appears due to local quantum and classical density fluctuations through the QCD droplet formation. Similar to our first point (see above), even at event-by-event analysis, the smoothing procedure is necessary to describe some stage of the collision process in the continuous medium approximation. We just smooth out the droplet picture for very high-density local fluctuations into an effectively hydrodynamic one.
@@
The realistic situation, however, can be more complicated, even at the selected criteria: in the case of low collision energies, the system may never reach complete thermalization, but only if possible, a partial one. In such a case, despite near full thermalization at the particlization stage at high enough energies, the system at the decaying (into the interacting particles) stage consists of two expanding components of baryon-rich matter: particle gas and fluid. The most natural way then is to select the energy density interval when the mechanism, forming the dense (QGP?) droplets, stops working, and soon after the corresponding (see above) hydrodynamical part of the system decays into hadrons without reaching full thermalization. In our approximation, it means that one has to transform the hydrodynamically involved part of the system into particles at this droplet’s (mean) decay time. So, because we do not know the quantum component of the QCD fluctuations, the corresponding energy density at which the decay of the hydro-component at small relativistic energies happens is a free parameter. The corresponding particle injection from this decay of the hydro-component in the case of not full thermalization is just added to the preserved yet UrQMD component. Of course, the local energy-momentum conservation law for the evolution and hydro-decays of the total system, consisting of both UrQMD (or SMASH) and hydro-components, must be implemented in any scenario. It is done in the iHKMe (like in iHKM) scenario that is developed in the presented article.
@@
However, as we already mentioned, switching from quasiclassical microscopic models (UrQMD, SMASH, etc) to the macroscopic hydrodynamic regime could be impossible if one naively tries to base it on a distribution function from a single transport simulation, as it brings significant fluctuations in coordinates and anisotropies in momentum space, contradicting the basic assumptions of hydrodynamics. So, one needs to use some averaging procedure to provide smooth initial conditions for a subsequent description of hydrodynamic expansion and, of course, especially, for event-by-event analysis.
@@
There are two common solutions to this problem. The straightforward one is to generate huge amounts of events and then average over them. However, to study the influence of fluctuations in the initial conditions on the final observables, one needs to introduce some similarity criteria between events. Of course, averaging over the centrality class might be too rough and unsuitable for event-by-event analysis. The common approach to this problem is applying some Gaussian smearing procedure to each particle and then constructing only the time component of the stress-energy tensor $T^{0\mu}$ and baryon current $J^{0}$. The other components are restored using an equation of state and explicit representation of the relativistic hydrodynamic tensors through macroscopic fields of velocity $u^{\mu}$, energy-density $\epsilon$, pressure $p$, and baryon density $n_{B}$.
@@
where $\bf u$} is the velocity of the particle, $\bf r$} is the vector from the particle’s position to the spatial point where its contribution is evaluated, and $\gamma$ is the Lorentz contraction factor. Such a procedure can be attributed to the averaging over an ensemble of “similar” collision events without generating them. The similarity between the events is described by the $\sigma$ parameter.
@@
Here, $r_{ij}^{\mu}=x^{\mu}_{i}-x^{\mu}_{j}$ represents the radius vector between the particle and the center of the cell $\left(x_{i},x_{j}\in\sigma^{\mu}\right)$, $n_{i}^{\mu}$ is the normal vector to hypersurface at point $x_{i}$, while $R$ is a free scalar parameter.
Many hydrodynamic models [40, 41, 42, 43], including the vHLLE code [44] employed in both the previous and current versions of the iHKM model, operate in Milne coordinates. These coordinates, defined by the proper time $\tau=\sqrt{t^{2}-z^{2}}$ and spacetime rapidity $\eta=\tanh^{-1}(z/t)$, are particularly suitable for high-energy collisions, where an approximately boost-invariant longitudinal expansion is expected. The use of hyperbolic (Milne) coordinates also allows one to fix the number of grid points in the longitudinal direction $N_{\eta}$ in advance, by scaling the grid spacing with proper time $\tau$ [45] instead of rescaling the full longitudinal extent of the system ($z_{\text{max}}$ grows almost with the speed of light), as would be necessary in Cartesian coordinates. This feature is especially useful for models that aim to describe a wide range of collision energies within a unified framework, as we do in iHKM.
As we use Milne coordinates for the system’s evolution in the subsequent stages, it is natural to choose a hypersurface of constant proper time $\tau$ in the kernel (8). For a fluid cell located at space-time rapidity $\eta$, the hypersurface element is given by
@@
where $\Delta x$, $\Delta y$, and $\Delta\eta$ are the cell sizes in the transverse and longitudinal directions, with numerical values of $0.3\,\mathrm{fm}$, $0.3\,\mathrm{fm}$, and $0.05$, respectively.
Utilizing kernel (8), we obtain the following inputs for the next stages of the model:
@@
Notably, we have not found any publications that apply particle smearing directly on a hypersurface of constant proper time $\tau$. Instead, such procedures are typically performed at constant laboratory time $t$, followed by a transformation to Milne coordinates. While we do not undertake a detailed comparison of these two approaches, we emphasize that the kernel (8) retains a Lorentz-covariant form, ensures particle number conservation, and accounts for Lorentz contraction along particle velocities (for more details, see Appendix A). Moreover, it enables an efficient reconstruction of the non-equilibrium distribution function on the particlization hypersurface (see Section II.4) by storing only 400 to 1000 particles (depending on the collision energy $\sqrt{s_{NN}}$) whose trajectories satisfy the selection criteria specified in Eq. (13)  [^1] With more common kernel 7 one needs to iterate over all tracks between $t_{\text{min}}=\tau$ and $t_{\text{max}}=\sqrt{\tau^{2}+z_{\text{max}}^{2}}$, where $z_{\text{max}}$ is longitudinal coordinate of leading particles at proper time $\tau$..
Lastly, we note that in the results presented in this paper, for simplicity, we do not consider separate conservation equations for electric charge and strangeness. Instead, we assume local constraints $n_{s}=0$ and $n_{q}=A/Z\cdot n_{B}\approx 0.4\,n_{B}$, appropriate for gold nuclei. Although both UrQMD and vHLLE support such calculations (but not the equations of state considered in this paper), we omit them here to reduce computational complexity and focus on the primary features of the model.
@@ II.2 Thermalization
The thermalization stage is one of the distinctive features of iHKM. During this stage, the matter can be phenomenologically decomposed into two distinct components: a (near) locally equilibrated component, described by macroscopic fields, and a non-thermalized component, represented by out-of-equilibrium hadrons and strings evolving via UrQMD in the cascade regime. Both components contribute to the non-equilibrium energy-momentum tensor [12] and the charge currents, ensuring that the corresponding equations respect the conservation laws for the total system. This stage is the primary difference between our model and more common models with instant thermalization
@@
where ${\cal P}_{\tau}={\cal P}(\tau)$ is a weight function such that ${\cal P}(\tau_{0})=1$ at the start of the thermalization stage, ${\cal P}(\tau_{th})=0$ at the end, and $0<{\cal P}(\tau_{0}<\tau<\tau_{th})<1$ in between. Its explicit form will be discussed later. Both the total stress-energy tensor and the baryon current obey conservation laws:
@@
$\partial_{\mu}T^{\mu\nu}_{\text{urqmd}}(x)=0,\quad\partial_{\mu}J_{\text{urqmd}}^{\mu}(x)=0\,$, (17)
@@
$\tilde{J}^{\mu}_{\text{hydro}}(x)=J^{\mu}_{\text{hydro}}(x)\cdot\left(1-{\cal P}_{\tau}\right),$ (21)